# SuperARC: Can Increasing Complexity Explain Intelligence? A Test for Artificial Super Intelligence Based On the Principles of Causal Recursive Compression and Algorithmic Probability

### v4

@@ 5.1 Next-digit Prediction Task with Binary and Non-binary Sequences
The objective of this experiment is to compare a fundamental characteristic of LLMs, that is, the prediction of the next token, with the power of understanding and then predicting approached through Algorithmic Probability Theory. This test, in particular, is inspired by  [77], which focuses on using Turing machines to approximate algorithmic complexity for short binary strings as a measure of algorithmic complexity as a means to explore fundamental principles of information and computational complexity, providing insights into the minimal description length of a string, an essential concept in understanding randomness and structured data.
@@
We tasked Large Language Models (LLMs) specializing in time series prediction with predicting the final digit of both non-binary sequences and binary sequences, the latter of which were categorised as either random or “climber" sequences. The results of the experiment involving binary sequences are presented in Figure 2.
@@
As shown in Figure 2, in the case of simple “climbers”, Lag-Llama achieved the best performance, with 70% precision, while TimeGPT–1 and Chronos barely reached 50% precision. However, for random sequences, which are considered highly complex, all models performed similarly, showing limited predictive power. This outcome suggests that, given the binary nature of the sequences, the models had a 50% chance of success, effectively reducing the task to guessing. These findings align with broader research that indicates that LLM models do not effectively capture sequential dependencies or complex patterns inherent in time series data. As highlighted by Tan et al. [78], despite their computational intensity, LLMs often fail to outperform simpler models, particularly when there is high complexity or randomness in the data.
@@
A comparable analysis was conducted using LLMs specialised in time-series data, using non-binary sequences of increasing complexity. In this test, a specific percentage of the final numbers in each sequence was required to be predicted. Three distinct metrics were utilised: general similarity, sort similarity, and the Levenshtein distance (refer to the section  4.4.2 for its definition). Figure 1 presents the results, where sort similarity and general similarity exhibit closely aligned trends. This indicates that the predictive accuracy of LLM models, even when fine-tuned for numerical series, diminishes as the complexity of the sequences increases. The resemblance between sort similarity and general similarity implies that while predictions may include some of the expected numbers, their correct order remains equally critical and may not always be achieved.
@@
Figure 2: Percentage of accuracy on binary climbers and random binary sequences by LLM models specialising in time series prediction compared with BDM. That climbers (up) where better predicted is expected from models that are able to intrinsically characterise and better predict simpler sequences. That TimeGPT performed better for random sequences than the other LLM models is a surprise.
@@
Figure 3 shows an increase in complexity as was expected, given the design of each group of generated sequences. The plot suggests that BDM can capture (and can generate) better complexity and randomness, since its values increase more consistently as complexity increases, unlike other measures. Shannon-entropy-based measures (and cognates) can account for statistical randomness only. Compression algorithms, for example, decrease as complexity increases, becoming more difficult to find regularities and increasing compression length as a function of complexity growth.
@@ 5.2 Free-form Generation Task with Non-binary Sequences
A subsequent analysis focused on the free-form test, where Large Language Models (LLMs) were given complete freedom to generate any model or formula capable of producing target sequences of increasing complexity.
Figure 10 shows the plots of complexity-related metrics for the models and formulas generated by LLMs used in this research. The metrics evaluated include the length of the LZW-compressed model, the length of the ZIP-compressed model, the BDM (Block Decomposition Method) of both the uncompressed model and its LZW and ZIP-compressed forms, and the Shannon entropy of the model.
@@ 5.2.1 Emergent abilities
Another experiment aimed to evaluate characteristics recently attributed to large language models (LLMs), particularly their so-called emergent abilities, which include innovation, discovery, and improvement. These attributes have been claimed to enable LLMs to perform at levels comparable to the human top 1% in fluency and originality, as suggested by Zhao et al. in their assessment of creativity in artificial intelligence systems [79].
@@ 5.3 Code Generation Task with Non-binary Sequences
A second test performed to evaluate compression was based on the no-compression percentage. According to this metric, a compressed–and therefore, comprehended–sequence could be expressed as a general (and ideally short) program. Print cases are considered here to have 100% non-compression, since they involve displaying the original sequence as is, which in our test is synonymous with not understanding the sequence.
Figure 4: Analysis of formulae generation for numerical sequences of increasing complexity. Top: Total number of valid generated formulae, where valid stands for different to ‘Not found’ response. Middle: Percentage of equivalence (output similarity among generated formulae) and accuracy (correct replication of target numeric sequences). Bottom: Distribution of formula types among accurate and total responses. The results highlight a direct correlation between sequence complexity and the model’s inability to generalise. Notably, the limitations of LLMs are particularly evident in contexts allowing complete freedom to find diverse yet correct solutions, underscoring an absence of creativity and genuine understanding, attributes often mistakenly attributed to these models [79].The newest version of ChatGPT–o1, Grok and Gemini performed worse than its preview version (see Sup. Inf).
Figure 5: Analysis of Python script generation for numerical sequences of increasing complexity. Top: Total number of scripts generated with valid results. Middle: Percentage of equivalence (output similarity among generated scripts) and accuracy (correct replication of target numeric sequences). Bottom: Distribution of script types among accurate and total responses. The findings challenge the presumed ability of LLMs to outperform humans in solving well-defined yet complex tasks. While high equivalence and some capacity for coherent solutions are observed at higher complexities, low accuracy highlights significant limitations. Despite extensive training in Python, the results confirm that without similar examples in the training dataset, it becomes extremely difficult–if not impossible–for LLMs to deduce solutions or generate multiple valid answers for the same problem. The newest version of ChatGPT–o1, Grok and Gemini performed worse than its preview version (see Sup. Inf).
@@
6 SuperARC-seq
Based on the previous experiments, it is possible to characterise one test directly related to the SuperARC framework: the SuperARC-seq. The objective of this test is to quantify intelligence and related cognitive capacities, specifically, reasoning and comprehension, drawing inspiration from the work in [77] and the theoretical and empirical studies here introduced. As mentioned, this test is grounded in one of the fundamental cognitive tasks: recognising patterns and evaluating the complexity of finite sequences, which inherently requires a level of understanding in order to provide a meaningful explanation. In our experiment, we generated short integer sequences (100 binary and 90 integer-valued in general, as seen in subsections 10.15 and 10.16, respectively, in Suppl. Information) and tasked several advanced LLMs with deriving a formula capable of reproducing the each of the target sequences.
@@ 6 SuperARC-seq
Thus, from these three types of correct results (i.e., the reconstructed sequence matches exactly the original one), we have four different classes of results: Correct & Non-Prints & Non-Ordinal; Correct & Ordinal; Correct & Prints; and Incorrect.
For any given tested model, the percentages of results belonging to each group can be combined as a vector of results, $\rho=[\%_{c,np,no},\%_{c,o},\%_{c,p},\%_{inc}]$, such that $\sum\rho_{i}=1$ as the percentages will be represented in the range [0,1] to resemble probabilities. We know, beforehand, that the best performing model would be one with $\rho_{best}=[1,0,0,0]$. Thus, a first possible test would be to check the overall percentage of correct answers.
@@
which would range from 0 to 1 for models that are not able to reproduce any sequence to models which perfectly reconstruct the sequences, respectively. However, this only accounts for the ability of LLMs to reproduce the initial sequence (planning) but not for their compression capabilities. To account for the latter, let us assume that the best possible algorithm for each element of the data set is $\mathcal{B}_{k,j}$, such that $\mathcal{B}_{k,j}()=D_{k,encoded}[j]$, and here the algorithm does not have a particular input, similar to the definition of algorithmic complexity. Thus:
$K(D_{k,encoded}[j])=K(\mathcal{B}_{k,j}())\leq K(\mathcal{B}_{k,j})$ (4)
due to the information non-increase theorem and to the fact that no inputs were used in the function. The ratio $K(D_{k,encoded}[j])/K(\mathcal{B}_{k,j})$ consistently falls within the range [0,1] for medium to long sequences when no embedding algorithms are employed. This behavior arises because approximations of algorithmic complexity are less reliable for short sequences, primarily due to the overhead inherent in theoretical computations. In order to surpass this limitation, since the difference between the true algorithmic complexity value and its approximation is bounded by a linear constant in general, instead of assessing the absolute algorithmic complexity (or any of its approximations), we shall consider a normalized version of it. To approximate algorithmic complexity, we will use the BDM/CTM approach, as described in detail in previous sections.
To build a normalized version of BDM/CTM, as pointed out in previous works [73], for any object of arbitrary size, it is possible to construct analogous objects that attain the minimum and maximum possible values of algorithmic complexity according to the Block Decomposition Method (BDM):
  • Minimum complexity object: This case is straightforward and corresponds to an object composed entirely of a single repeating symbol - for instance, a binary string consisting solely of zeros.
@@
The primary advantage of considering a normalized measure lies in its ability to enable comparisons between objects of varying sizes, effectively mitigating the influence of size on the measure itself. This property is particularly in the case of the present study, where we compare complexities of sequences and formulas generating them.
This way, the following ratio presents itself as an interesting weighting factor for the probabilities in equation 3:
@@
The ratio in equation 5 measures how the algorithmic complexity of the formula and sequence compare to the other possible outputs of the LLM. If relative algorithmic complexity (measured by the normalized BDM value) for the formula is greater than it was for the sequence itself, this suggests the LLM did not success in compressing the input sequence (it made the formula have a greater relative algorithmic complexity). On the other hand, if the opposite occurs, then the LLM could compress the sequence comparatively to other possible outputs of the LLM. The ratio in equation 5 ranges from 0 to a positive value $M>1$, which happens when the best possible compression is achieved (the inverse mapping of CTM). Since $M$ is not known beforehand, we can use a nonlinear mapping that saturates the value of the ratio to a maximum value of 1 (similar to an activation function). The hyperbolic tangent function can be used in this case, since $\text{tahn}(0)=0$ and $\lim_{x\to\infty}\text{tahn}(x)=1$. Thus, a candidate weighting factor for the probabilities in 3 is:
$\delta_{k,j}=\text{tanh}\left(\frac{nBDM(D_{k,encoded}[j])}{nBDM(\mathcal{B}_{k,j})}\right)$ (6)
@@
It can be seen that $\varphi\in[0,1]$ encompasses different behaviours. For example, $\varphi\in[0,0.01]$ if only print-type models are outputted. Also, $\varphi\in[0,0.1]$ if only ordinal-like formulas are created. Finally, $\varphi\in[0,1]$ in cases where the LLMs create formulas that are always correct, do not copy nor create ordinal mappings. The ranges will be populated with varying compression levels corresponding to the algorithms obtained. Overall, if the score is 0, all the formulas were wrong. If it is 0.5, it can represent the case where half the outputs were correct and half wrong, with the formulas produced with highest compression levels. So, in a regular half and half case, since compression will not be optimal, the test score is less than 0.5.The test performance results for each model are calculated using equation 9 for $\mathcal{T}$ in Algorithm 1.
There are some possible variations for the test metric in equation 9. For example, some sort of Bayesian approach could be used to consider that the elements of $\rho$ are not constants, but random variables which could account for the number of different correct/incorrect answers for the same input sequence. In this way, the multiplicity of possible generators is taken into account, better capturing the concept of algorithmic probability, and the output of the test would be a random variable instead. However, LLMs hardly produced even one correct answer, therefore we kept the formula simple.
As described, equation 9 tests for two features, compression via non-print computer programs and non-ordinal mathematical formulas to the input sequence, and prediction, by running all programs and all formulas to match each sequence digit, and penalising them when they did not represent an actual compressed model that generated a possible new digit of the sequence when run in reverse, i.e. when ‘decompressed’. The test formula assigns greater importance to correct cases that are not solutions of the type ‘print($s$)’ where $s$ is the sequence for which the AI system is asked for a model, given that a print model does not allow generalisation by prediction through simulation, as running a print command will only print up to the last digit. The same is true for what we call ‘ordinals’, which is simply indicating the index of the non-zero non-one element in the binary (or binary embedded) sequence, meaning that, together with the ‘print’ case, the system failed in its attempts at abstracting features of the object. Finally, the formula punishes ordinal and print answers in a weighted fashion. The best performer can only reach a $\varphi$ of 1 while the lowest value is 0.
6.1 Applying SuperARC-seq

### v5

@@ 2 Results
The tests were inspired by, and based on, two methods called Coding Theorem and Block Decomposition methods (CTM and BDM) [33, 34, 35], which use a composition of pattern-matching and running a very large set of small computer programs to approximate a distribution allowing the estimation of the algorithmic complexity of short objects providing insights into the minimal description length of objects such as strings, sequences or images.
These methods have been applied to humans before on similar tasks and the same tools (CTM/BDM), showing that the methods can capture aspects that other intelligence tests based on pattern matching fail or require much more ad hoc information and assumptions to reproduce them [2]. Humans showed some predictive capabilities that could be quantified only with CTM/BDM and not with other statistical tools. This led CTM and BDM to be widely used today in the psychometric testing space by multiple groups [36, 37, 38].
Among the results in [2], people showed that at 25 years of age they had the highest ability to identify and produce the highest random complexity when required and decreased when older. More experiments should be conducted to test human’s recursive prediction capabilities, but what the article showed [2] was that people built and had better perception models for producing and identifying random versus non-random content when they were 25 than at any other age and therefore were more sensitive to identifying less predictive data, all of which are compatible with current knowledge of human cognitive trends and capabilities.
While we do not necessarily expect humans to perform well on predictive tests such as those introduced here, humans are, in principle, mechanistically capable of solving them. As such, these tasks are arguably within the definition of AGI as a system possessing the full range of human capabilities, without time constraints. Moreover, in the context of ASI, the natural assumption is that of optimal prediction, for which Algorithmic Information Theory (AIT)–used and exploited here for testing purposes–is the accepted mathematical framework.
@@ 2.1 Next-digit Prediction Task with Binary and Non-binary Sequences
The objective of this experiment is to test a fundamental property of LLMs, that is, the prediction of the next token, within the context of Algorithmic Information Theory (AIT), the accepted mathematical theory governing optimal prediction, as introduced in Section 7.1.
@@
We tasked Large Language Models (LLMs) specializing in time series prediction with predicting the final digit of both non-binary sequences and binary sequences, the latter of which were categorised as either random or “climber” sequences. The results of the experiment involving binary sequences are presented in Figure 1.
We call “climbers” those sequences that have some recursive properties that make them stand out and have been properly ranked as of lower complexity given their recursive properties according to CTM/BDM (see Section 7.12.10). As shown in Figure 1, in the case of what we call “climbers”, Lag-Llama achieved the best performance, with 70% precision, while TimeGPT—1 and Chronos barely reached 50% precision, while CTM/BDM was used as a gold standard.
@@
However, for random sequences, which are considered highly ‘complex’ in this context, all models performed similarly, showing limited predictive power, as expected. These results suggest that, given the binary nature of the sequences, the models had a 50% chance of predictive success, effectively reducing the task to guessing, yet some performance difference was noticed for non random cases, indicating some predictive ability. However, their performance aligns with broader research that indicates that LLM models do not effectively capture sequential dependencies or complex patterns inherent in time series data. As highlighted by Tan et al. [39], despite their computational intensity, LLMs often fail to outperform simpler models, particularly when there is high complexity or randomness in the data.
@@
A comparable analysis was conducted using LLMs specialised in time-series data, using non-binary sequences of increasing complexity. In this test, a specific percentage of the final numbers in each sequence was required to be predicted. Three distinct metrics were utilised: general similarity, sort similarity, and the Levenshtein distance (refer to Section 6.7.2 for its definition). Figure 9 presents the results, where sort similarity and general similarity exhibit closely aligned trends. This indicates that the predictive accuracy of LLM models, even when fine-tuned for numerical series, diminishes as the complexity of the sequences increases. The resemblance between sort similarity and general similarity implies that while predictions may include some of the expected numbers, their correct order remains equally critical and may not always be achieved.
@@
Figure 1: Percentage of accuracy on binary “climbers” and random binary sequences by LLM models specialising in time series prediction compared with BDM. That climbers (up) were better predicted is expected from models that are able to intrinsically characterise and better predict simpler sequences. Sequence prediction is a fundamental problem in science, from genetics to protein folding in biology to digital twin technology in medicine and healthcare.
@@
Figure 2 shows an increase in complexity as expected, given the design of each group of generated sequences. The plot suggests that BDM can capture (and can generate) better complexity and randomness, since its values increase more consistently as complexity increases, unlike other measures. Shannon entropy-based measures (and cognates) can account for statistical randomness only. Compression algorithms, for example, decrease as complexity increases, becoming more difficult to find regularities and increasing compression length as a function of complexity growth.
@@ 2.2 Free-form generation task with non-binary sequences
A subsequent analysis focused on the free-form test, where LLMs were given complete freedom to generate any model or formula capable of producing target sequences of increasing complexity.
Figure 10 shows the plots of complexity-related metrics for the models and formulas generated by LLMs used in this research. The metrics evaluated include the length of the LZW-compressed model, the length of the ZIP-compressed model, the BDM of both the uncompressed model and its LZW and ZIP-compressed forms, and the Shannon entropy of the model.
@@ 2.2.1 Emergent abilities
Another experiment aimed to evaluate characteristics recently attributed to LLMs, particularly their so-called emergent abilities, which include innovation, discovery, and improvement. (See also Sections 6.2 and 7.2). These attributes have been claimed to enable LLMs to perform at levels comparable to the human top 1% in fluency and originality, as suggested by Zhao et al. in their assessment of creativity in artificial intelligence systems [40].
@@ 2.3 Code generation task with non-binary sequences
A second test performed to evaluate compression was based on the no-compression percentage. According to this metric, a compressed—and therefore, comprehended—sequence could be expressed as a general (and ideally short) program. Print cases are considered here to have 100% non-compression, since they involve displaying the original sequence as is, which in our test is synonymous with not understanding the sequence.
Figure 3: Comprehensive analysis of formulae generation for numerical sequences of increasing complexity. Top left: Percentage of equivalence between generated formulae, measuring output similarity and solution diversity. Top right: Accuracy rates showing correct replication of target numeric sequences across complexity levels. Bottom: Integrated view combining formula generation volume (gold line, secondary axis) with type distribution among both total (lighter bars) and accurate (darker bars) responses, categorised as known sequences (blue), pure mathematical expressions (green), and not found (red). The results demonstrate a direct correlation between sequence complexity and diminished model performance, with particularly stark degradation in equivalence rates suggesting limited solution diversity. The integrated bottom panel reveals that whilst models may generate valid formulae at lower complexities, the proportion of accurate responses declines precipitously, and reliance on known sequences dominates over novel mathematical reasoning. These limitations are especially pronounced in contexts permitting complete freedom to discover diverse yet correct solutions, underscoring an absence of genuine creativity and mathematical understanding, attributes often mistakenly attributed to these models [40]. Notably, newer versions of ChatGPT-o1, Grok, and Gemini performed worse than their preview iterations (see Supplementary Information).
Figure 4: Comprehensive analysis of language model performance in Python script generation across complexity levels (Low, Medium, High). Up: Equivalence percentage (left) and accuracy (right) versus complexity. Bottom: For each model, semi-transparent left bars show total script type distribution (Known sequence=red, Not found=blue, Pure math=green, Print=orange); solid right bars show accurate predictions only; gold diamonds (right y-axis) indicate valid script volume. Disparity between left/right bar heights quantifies the accuracy gap. Results expose fundamental LLM limitations: whilst models generate coherent solutions, accuracy deteriorates markedly with complexity. Predominance of ‘Not found’ (blue) at higher complexities indicates systematic failure to recognise solution strategies. Upper trajectories show equivalence remains stable whilst accuracy plummets—models generate internally consistent but incorrect approaches. Without analogous training exemplars, LLMs cannot reliably deduce solutions despite extensive Python training. Notably, newer iterations (ChatGPT-5, Grok, Gemini) underperformed preview versions (see Supplementary Information), challenging assumptions of monotonic improvement.
@@
2.4 SuperARC-seq
Based on the previous experiments, it is possible to characterise one test directly related to the SuperARC framework: the SuperARC-seq. The objective of this test is to quantify intelligence and related cognitive capacities, specifically, reasoning and comprehension, drawing inspiration from the work in [34] and the theoretical and empirical studies here introduced. As mentioned, this test is grounded in one of fundamental cognitive tasks: recognising patterns and evaluating the complexity of finite sequences, which inherently requires a level of comprehension in order to provide a meaningful explanation. In our experiment, we generated short integer sequences (100 binary and 90 integer-valued in general, as seen in subsections 7.12.11 and 7.12.12, respectively, in Sup. Inf.) and tasked several advanced LLMs with deriving a formula capable of reproducing each of the target sequences.
@@ 2.4 SuperARC-seq
Thus, from these three types of correct results (i.e., the reconstructed sequence matches exactly the original one), we have four different classes of results: Correct & Non-Both; Correct & Ordinal; Correct & Prints; and Incorrect.
For any given tested model, the percentages of results belonging to each group can be combined as a vector $\rho$ of rates in the range $\left[0,1\right]$, where $\rho=[\%_{c,np,no},\%_{c,o},\%_{c,p},\%_{inc}]$ such that $\sum_{i=1}^{4}\rho_{i}=1$. Notice that the percentages are represented in the range $\left[0,1\right]$ in order to resemble probabilities. We know, beforehand, that the best-performing model would be one with $\rho_{best}=[1,0,0,0]$. Thus, a first possible test would be to check the overall percentage of correct answers.
@@
which would range from 0 to 1 for models that are not able to reproduce any sequence to models which perfectly reconstruct the sequences, respectively. However, this only accounts for the ability of LLMs to reproduce the initial sequence (planning) but not for their compression capabilities. To account for the latter, let us assume that the best possible algorithm for each element of the data set is $\mathcal{B}_{k,j}$, such that $\mathcal{B}_{k,j}()=D_{k,encoded}[j]$, and here the algorithm does not have a particular input, similar to the definition of algorithmic complexity. Thus, from the basic properties in AIT:
$K(D_{k,encoded}[j])=K(\mathcal{B}_{k,j}())\leq K(\mathcal{B}_{k,j})+\mathbf{O}(1)$ (2)
The ratio
$\frac{K\left(D_{k,encoded}[j]\right)}{K(\mathcal{B}_{k,j})+\mathbf{O}(1)}$ (3)
consistently falls within the range [0,1] for medium to long sequences when no embedding algorithms are employed. This behaviour arises because approximations of algorithmic complexity are less reliable for short sequences, primarily due to the overhead inherent in theoretical computations. In order to surpass this limitation, instead of assessing the absolute algorithmic complexity (or any of its approximations), we shall consider a normalised version of it (denoted by $nBDM(\cdot)$).
To approximate algorithmic complexity, we will use the BDM/CTM approach, as described in detail in previous sections, and its normalised version, as pointed out in previous works [35] for any object of arbitrary size, it is possible to construct analogous objects that attain the minimum and maximum possible values of algorithmic complexity according to the BDM:
  • minimum complexity object: This case is straightforward and corresponds to an object composed entirely of a single repeating symbol—for instance, a binary string consisting solely of zeros;
@@
The primary advantage of considering a normalised measure lies in its ability to enable comparisons between objects of varying sizes, effectively mitigating the influence of size on the measure itself. This property is particularly in the case of the present study, where we compare complexities of sequences and formulas generating them.
This way, the following ratio presents itself as an interesting weighting factor for the probabilities in Equation (1):
@@
The ratio in Equation (4) measures how the algorithmic complexity of the formula and sequence compare to the other possible outputs of the LLM. If the relative algorithmic complexity (measured by the normalised BDM value) of the formula is greater than it was for the sequence itself, this suggests the LLM did not success in compressing the input sequence (it made the formula have a greater relative algorithmic complexity). On the other hand, if the opposite occurs, then the LLM could compress the sequence comparatively to other possible outputs of the LLM.
The ratio in Equation (4) ranges from 0 to a positive value $M>1$, which happens when the best possible compression is achieved (the inverse mapping of CTM). Since $M$ is not known beforehand, we can use a nonlinear mapping that saturates the value of the ratio to a maximum value of 1 (similar to an activation function). The hyperbolic tangent function can be used in this case, since $\tanh(0)=0$ and $\lim_{x\to\infty}\tanh(x)=1$. Thus, a candidate weighting factor for the probabilities in 1 is:
$\delta_{k,j}=\tanh\left(\frac{nBDM(D_{k,encoded}[j])}{nBDM(\mathcal{B}_{k,j})}\right)$ (5)
@@
It can be seen that $\varphi\in[0,1]$ encompasses different behaviours. For example, $\varphi\in[0,0.01]$ if only print-type models are outputted. Also, $\varphi\in[0,0.1]$ if only ordinal-like formulas are created. Finally, $\varphi\in[0,1]$ in cases where the LLMs create formulas that are always correct, do not copy nor create ordinal mappings. The ranges will be populated with varying compression levels corresponding to the algorithms obtained. Overall, if the score is 0, all the formulas were wrong. If it is 0.5, it can represent the case where half the outputs were correct and half wrong, with the formulas produced with highest compression levels. So, in a regular half and half case, since compression will not be optimal, the test score is less than 0.5. The test performance results for each model are calculated using Equation (8) for $\mathcal{T}$ in Algorithm 1.
There are some possible variations for the test metric in Equation (8). For example, some sort of Bayesian approach could be used to consider that the elements of $\rho$ are not constants, but random variables which could account for the number of different correct/incorrect answers for the same input sequence. In this way, the multiplicity of possible generators is taken into account, better capturing the concept of algorithmic probability, and the output of the test would be a random variable instead. However, LLMs hardly produced even one correct answer, therefore we kept the formula simple.
As described, Equation (8) tests for two features, compression via non-print computer programs and non-ordinal mathematical formulas to the input sequence, and prediction, by running all programs and all formulas to match each sequence digit, and penalising them when they did not represent an actual compressed model that generated a possible new digit of the sequence when run in reverse, i.e. when ‘decompressed’. The test formula assigns greater importance to correct cases that are not solutions of the type ‘print($s$)’ where $s$ is the sequence for which the AI system is asked for a model, given that a print model does not allow generalisation by prediction through simulation, as running a print command will only print up to the last digit. The same is true for what we call ‘ordinal’, which is simply indicating the index of the non-zero non-one element in the binary (or binary embedded) sequence, meaning that, together with the ‘print’ case, the system failed in its attempts at abstracting features of the object. Finally, the formula punishes ordinal and print answers in a weighted fashion. The best performer can only reach a $\varphi$ of 1 while the lowest value is 0.
2.4.1 Applying SuperARC-seq